General Informations

Status-Key

Status/green.gif

Resolved

Status/orange.gif

Still working but with some errors

Status/red.gif

Pending


Maintenance Reboot of Solaris Server Yosemite

STATUS: Status/green.gif

2010-12-07: 7:30 AM
Server yosemite has been rebooted successfully. All services are available.
2010-12-07: 7:00 AM
Due to a shortage of available memory we are forced to reboot the solaris server yosemite. Downtime approx. 30 minutes.


cooling water system outage on clusters

STATUS: Status/green.gif

2010-11-26: 5:00 PM
Host autserv02 is running as well. All hosts can be used.
2010-11-26: 4:40 PM
Server racks are cooled again, all hosts except of autserv02 are running and can be used.
2010-11-26: 4:00 PM

Server racks are still down. --> Update follows at 5 PM or earlier

2010-11-26: 3:10 PM

One of the cooling water pumps installed in ETZ/D/96.2 does not work correctly. This forces some of the racks in this server room to shutdown in order to protect the servers from thermal damage. clusters from IFH, IBT, BIWI, TIK, IKT and VAW are affected. the facility management is working on solving the problem. --> Update follows at 4 PM


email phishing attack

STATUS: Status/green.gif

2010-11-23

Yesterday between 18:20 and 19:50 about 320 Phishing Mails have been sent to different Users at D-ITET. The Mails pretend to come from IT Support Group and contain the subject ISG.EE Webmail Alert. The mail tells something about spammers that have compromised the ISG.EE Webmail Account and that you should provide your Username, Password and some Alternate Email. Please remember, that the ISG.EE Team will NEVER ask you for your Password! If you still have replied to this phishing mail please contact us immediately under support@ee.ethz.ch so that we can plan with you the next steps to keep your account safe.


oenone home server crash

STATUS: Status/green.gif

2010-11-17

During this night at around 00:15 oenone one of our home-servers crashed. Users with homes on oenone where affected, these are BIWI, VAW, Collegium Helveticum, Control, IBT, IKT. The server is now checking the filesystems and comming up again.

We are sorry for the caused inconvenience and we are investigating the problem.

Update: 08:00
oenone is now up and running again.
Update: 2010-11-18 07:30
We opened a support case at Sun/Oracle for this server.


cooling water system outage for some clusters

STATUS: Status/green.gif

2010-11-16

On last friday evening one of the cooling water pumps installed in ETZ/D/96.2 stopped working correctly. This forced some of the racks in this server room to shutdown in order to protect the servers from thermal damage. All clusters from IFH, IBT, BIWI, TIK, IKT and VAW were affected.

The facility management is working on solving the problem.

The servers are currently (08:35) down again. Please, even if they come up again, do not use them for long-timed computations as we still do not know when exactly the technician has solved the issue.

Update: 2010-11-17 08:25
The rack systems are running now with only one cooling water pump. A new pump is ordered by the rack company.
Update: 2010-11-18 16:00
Planed substitution of broken pump will be on 25.11 or 26.11.


CategoryEDUC

Status/Archive/2010 (last edited 2016-01-06 12:43:03 by alders)